# Spin-resolved quasiparticle interference patterns on altermagnets via non-spin-resolved scanning tunneling microscopy

### v1

@@ I Introduction
A rather simple model for realizing altermagnetic order is the two-dimensional Lieb lattice [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The Lieb lattice has a four-site unit cell in which only three sites are occupied by atoms, with the two diagonally opposed ones possessing a magnetic moment and the last one being non-magnetic. This lattice symmetry is invariant under a fourfold rotation followed by time reversal, but it breaks the usual translation symmetry that would interchange the two magnetic sublattices of a Néel antiferromagnet. As a result, the Lieb lattice can support an altermagnetic state.
Finding a stable 2D electronic Lieb lattice has been challenging [32, 33]. The original motivation for introducing the Lieb lattice was to describe layers in high $T_{c}$ superconductors [21]. Meanwhile, several recently proposed materials have Lieb-like layers and are altermagnetic candidates. For example, the layered oxychalcogenide KV2Se2O crystallizes in a structure that within the layers realizes a 2D Lieb lattice of vanadium and oxygen atoms. It has been demonstrated to be a metallic, room-temperature altermagnet with a $d$-wave spin-momentum locking of its Fermi surface [34]. The related compound Rb1-δV2Te2O is another layered antiferromagnet now identified as a room-temperature altermagnet exhibiting $C$-paired spin-valley locking consistent with altermagnetic order [35]. Likewise, the Mn-based insulator La2O3Mn2Se2 consists of Mn2O layers that form a Lieb lattice. It has been shown to host an antiparallel order with the expected altermagnetic band splitting [36, 37]. These examples underscore that Lieb-like sublattice geometries (and related structures) provide fertile ground for altermagnetism.
@@ II Model
Figure 1: (a) Two-dimensional altermagnetic crystal lattice with two sites hosting opposite magnetic moments (up/red and down/blue) and two non-magnetic sites (white) in a checkerboard pattern. Itinerant electrons couple to the local moments through an exchange term $\pm J\sigma$, with $\sigma=\pm 1$ for spin-up and spin-down carriers respectively. The values $\epsilon_{2/3}$ represent on-site energies for the non-magnetic sites and $t$ the hopping strength between nearest-neighbor bonds. By tuning $\epsilon_{2}$ and $\epsilon_{3}$, the system can be tuned between an antiferromagnetic (AFM) state with $\epsilon_{2}=\epsilon_{3}$ and an altermagnetic (AM) state with $\epsilon_{2}\neq\epsilon_{3}$. (b) Band structure of the model along the high-symmetry-points for the parameters $(\epsilon_{2}=\epsilon_{2}=1)$ in black and $(\epsilon_{2}=0,\epsilon_{2}=1)$ in red and blue. In the case of $\epsilon_{2}\neq\epsilon_{3}$, the altermagnetic $d$-wave spin splitting becomes visible along the $\Gamma\to X$ and $\Gamma\to Y$ paths. (c) Fermi surface of the model.
@@
$\mathcal{H}=$ $-t\sum_{\langle i,j\rangle,\sigma}c_{i,\sigma}^{\dagger}c_{j,\sigma}-J\sum_{i,\sigma}\sigma S_{i}^{z}c_{i,\sigma}^{\dagger}c_{i,\sigma}$
@@
Here, $c_{i,\sigma}^{\dagger}$ and $c_{i,\sigma}$ denote the creation and annihilation of an electron with spin $\sigma$ at lattice site $i$, respectively, while $t$ denotes the hopping strength between two nearest neighbors. A Kondo-like s-d coupling between the magnetic moment of the itinerant electrons and the magnetic moment of the local magnetic sites is added by the second term with a coupling strength $J$. For simplicity, we focus on a coupling along the $z$ direction, which is sufficient to observe altermagnetism in this model. We absorb the magnitude of $S_{i}^{z}$ into $J$ and assume site 1 has a local spin up and site 4 has local spin down. Sites 2 and 3 are nonmagnetic and have on-site energies $\epsilon_{\text{NM}}\in(\epsilon_{2},\epsilon_{3})$. The Fermi level is controlled by $\mu$.
@@ III Local density of states
In this section, we analyze the real space local density of states (LDOS) to map out how electrons distribute across the lattice. For this purpose, we dicsretize our tight-binding Hamiltonian (1) on an $N\times N$ grid and assemble it as a $2N^{2}\times 2N^{2}$ matrix $H$, so that $\mathcal{H}=\Psi^{\dagger}H\Psi$ with
@@
The total electron occupation at each site is then $\rho_{i}(\omega)=\sum_{\sigma}\rho_{i,\sigma}(\omega)$. The LDOS is calculated at the Fermi level ($\omega=0$) with $\mu$ chosen in a way that the Fermi level crosses the lowest two bands, regardless of how $\epsilon_{1}$ and $\epsilon_{2}$ may shift the bands. For the four-site and Lieb lattice altermagnetic cases we choose $\epsilon_{3}>\epsilon_{2}$. We also calculate the spin resolved effect of the impurity as $\Delta\rho_{i,\sigma}=\rho_{i,\sigma,\text{impure}}-\rho_{i,\sigma,\text{clean}}$ in order to explicitly show how the impurity affects the two spin flavors differently. We set the lattice constant to be $a=1$ in all further calculations following this section. All other system parameters are noted in Table 1.
@@ Acknowledgements.
We thank Changan Li and Johannes Mitscherling for useful discussions. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project SFB 1170 and DFG through the Würzburg-Dresden Cluster of Excellence ct.qmat (EXC 2147, project-id 390858490).

### v2

@@ I Introduction
A rather simple model for realizing altermagnetic order is the two-dimensional Lieb lattice [24, 5, 28, 21, 10, 17, 2, 23, 26, 41, 20, 42, 11, 45]. The Lieb lattice has a four-site unit cell in which only three sites are occupied by atoms, with the two diagonally opposed ones possessing a magnetic moment and the last one being non-magnetic. This lattice symmetry is invariant under a fourfold rotation followed by time reversal, but it breaks the usual translation symmetry that would interchange the two magnetic sublattices of a Néel antiferromagnet. As a result, the Lieb lattice can support an altermagnetic state.
Finding a stable 2D electronic Lieb lattice has been challenging [35, 44]. The original motivation for introducing the Lieb lattice was to describe layers in high $T_{c}$ superconductors [24]. Meanwhile, several recently proposed materials have Lieb-like layers and are altermagnetic candidates. For example, the layered oxychalcogenide KV2Se2O crystallizes in a structure that within the layers realizes a 2D Lieb lattice of vanadium and oxygen atoms. It has been demonstrated to be a metallic, room-temperature altermagnet with a $d$-wave spin-momentum locking of its Fermi surface [19]. The related compound Rb1-δV2Te2O is another layered antiferromagnet now identified as a room-temperature altermagnet exhibiting $C$-paired spin-valley locking consistent with altermagnetic order [48]. Likewise, the Mn-based insulator La2O3Mn2Se2 consists of Mn2O layers that form a Lieb lattice. It has been shown to host an antiparallel order with the expected altermagnetic band splitting [43, 13]. See also Ref. [6] for a first-principle study providing further material candidates. These examples underscore that Lieb-like sublattice geometries (and related structures) provide fertile ground for altermagnetism.
@@ II Model
Figure 1: (a) Two-dimensional altermagnetic crystal lattice with two sites hosting opposite magnetic moments (up/red and down/blue) and two non-magnetic sites (white) in a checkerboard pattern. Itinerant electrons couple to the local moments through an exchange term $\pm J\sigma$, with $\sigma=\pm 1$ for spin-up and spin-down carriers respectively. The values $\epsilon_{2/3}$ represent on-site energies for the non-magnetic sites and $t$ the hopping strength between nearest-neighbor bonds. By tuning $\epsilon_{2}$ and $\epsilon_{3}$, the system can be tuned between an antiferromagnetic (AFM) state with $\epsilon_{2}=\epsilon_{3}$ and an altermagnetic (AM) state with $\epsilon_{2}\neq\epsilon_{3}$. (b) Band structure of the model along the high-symmetry-points for the parameters $(\epsilon_{2}=\epsilon_{3}=1)$ in black and $(\epsilon_{2}=0,\epsilon_{3}=1)$ and $(\epsilon_{2}=1,\epsilon_{3}\to\infty)$ in red (spin up) and blue (spin down). In the case of $\epsilon_{2}\neq\epsilon_{3}$, the altermagnetic $d$-wave spin splitting becomes visible along the $\Gamma\to X$ and $\Gamma\to Y$ paths. (c) Fermi surface of the model.
@@
$\mathcal{H}=$ $-t\sum_{\langle i,j\rangle,\sigma}c_{i,\sigma}^{\dagger}c_{j,\sigma}-J\sum_{i,\sigma\sigma^{\prime}}\mathbf{S}_{i}\cdot c_{i,\sigma}^{\dagger}\boldsymbol{\sigma}_{\sigma\sigma^{\prime}}c_{i,\sigma^{\prime}}$
@@
Here, $c_{i,\sigma}^{\dagger}$ and $c_{i,\sigma}$ denote the creation and annihilation of an electron with spin $\sigma$ at lattice site $i$, respectively, while $t$ denotes the hopping strength between two nearest neighbors. A Kondo-like s-d coupling between the magnetic moment of the itinerant electrons and the magnetic moment of the local magnetic sites is added by the second term with a coupling strength $J$ in Eq. 1. $\boldsymbol{\sigma}$ is a vector of Pauli matrices, such that $\sum_{\sigma\sigma^{\prime}}c_{i,\sigma}^{\dagger}\boldsymbol{\sigma}_{\sigma\sigma^{\prime}}c_{i,\sigma^{\prime}}$ represents the spin of the itinerant electron at site $i$. We separate the itinerant electrons and localized spins as two separate degrees of freedom and focus on Néel type ordering of the localized spins $\mathbf{S}_{i}$. Without loss of generality, we assume the localized spins order along the $z$-direction, for instance, due to easy-axis anisotropy. Then, the s-d coupling term becomes $J\sum_{i,\sigma}\sigma S_{i}^{z}c_{i,\sigma}^{\dagger}c_{i,\sigma}$. We absorb the magnitude of $S_{i}^{z}$ into $J$ and assume site 1 has a local spin up and site 4 has local spin down. Sites 2 and 3 are nonmagnetic and have on-site energies $\epsilon_{\text{NM}}\in(\epsilon_{2},\epsilon_{3})$. The Fermi level is controlled by $\mu$. We assume the effect of spin-orbit coupling on the electron band structure is negligible compared to the non-relativistic spin splitting, as expected in, e.g., KV2Se2O [19].
@@ III Local density of states
In this section, we analyze the real space local density of states (LDOS) to map out how electrons distribute across the lattice. For this purpose, we discretize our tight-binding Hamiltonian (1) on an $N\times N$ grid with hard-wall boundary conditions and assemble it as a $2N^{2}\times 2N^{2}$ matrix $H$, so that $\mathcal{H}=\Psi^{\dagger}H\Psi$ with
@@
The total electron occupation at each site is then $\rho_{i}(\omega)=\sum_{\sigma}\rho_{i,\sigma}(\omega)$. The LDOS is calculated at the Fermi level ($\omega=0$) with $\mu$ chosen in a way that the Fermi level crosses the lowest two bands, regardless of how $\epsilon_{1}$ and $\epsilon_{2}$ may shift the bands. For the four-site and Lieb lattice altermagnetic cases we choose $\epsilon_{3}>\epsilon_{2}$. We also calculate the spin resolved effect of the impurity as $\Delta\rho_{i,\sigma}=\rho_{i,\sigma,\text{impure}}-\rho_{i,\sigma,\text{clean}}$ in order to explicitly show how the impurity affects the two spin flavors differently. We set the lattice constant to be $a=1$ in all further calculations following this section. All other system parameters are noted in Table 1. The LDOS can be directly measured in STM experiments, if the tip is carefully chosen such that its density of states is flat in the relevant energy range.
@@ IV Quasiparticle interference pattern
A potential challenge in experiments is how to determine which magnetic sublattice the impurity is placed on. STM topography measurements can be used to find the location of the impurity, and assuming a Lieb like lattice, the two magnetic sites can be identified without any spin-resolved STM measurement (assuming the magnetic order has been confirmed by other means). Then, one does not need to identify which sublattice is spin up and which is spin down, simply that two measurements done on the same geometry have the impurity placed on two different magnetic sublattices. The predicted difference in the QPI patterns for the two cases can then be checked, and the results can be interpreted as the two spin-split Fermi surfaces, without necessarily knowing which is spin up and which is spin down. We also perform calculations with several impurities placed on the same sublattice and find similar QPI patterns.
@@ Acknowledgements.
We thank Changan Li, Johannes Mitscherling, Artem Odobesko, and Pavlo Sukhachov for useful discussions. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project SFB 1170 and DFG through the Würzburg-Dresden Cluster of Excellence ct.qmat (EXC 2147, project-id 390858490).